OpenStudy (anonymous):

Solve the open sentence. –1 ≤ n + 2 ≤ 6 A.n ≤ –3 and n ≤ 2 B.n ≥ –3 and n ≤ 4 C.n ≥ 3 and n ≤ –2 D.n ≤ 1 and n ≤ 6

OpenStudy (anonymous):

@Willie579

OpenStudy (anonymous):

@Preetha @Hoslos @ganeshie8 PLZ HELP

OpenStudy (lochana):

okay. our inequality is –1 ≤ n + 2 ≤ 6 we can make this into two parts \[-1≤ n +2 \ \ and \ \ n+2 ≤ 6\]

OpenStudy (lochana):

okay?

OpenStudy (anonymous):

got u

OpenStudy (anonymous):

how u get 6

OpenStudy (lochana):

so now take first one it is \[-1 ≤ n +2\]\[-1-2 ≤n +2 -2 \]\[-3≤ n\]

OpenStudy (lochana):

it is the same as saying \[n \ge -3\]

OpenStudy (lochana):

now take second part it is \[n+2 \le 6\]\[n + 2 -2 \le 6 -2\]\[n \le 4\]

OpenStudy (lochana):

answer is B

OpenStudy (anonymous):

thx for making me understand

OpenStudy (lochana):

you are welcome. bye:)

OpenStudy (anonymous):

see ya